OpenStudy (anonymous):

find the integral of 1/square root of 4-x^2 dx

OpenStudy (anonymous):

you are supposed to make a trig sub, but you can do it without, if you recall that \[\frac{d}{dx}\sin^{-1}(x)=\frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

\[\int\limits \frac{ 1 }{ \sqrt{a^2-x^2}}dx=\sin^{-1} \left( \frac{ x }{a } \right)\]

OpenStudy (anonymous):

or just go right to the answer

OpenStudy (anonymous):

I dont get it. I'm a visual person, I need a step by step.

OpenStudy (anonymous):

@surjithayer gave you the final form, something that is probably in the back cover of your text

OpenStudy (anonymous):

or you can rewrite \[\frac{1}{\sqrt{4-x^2}}\] as \[\frac{1}{2}\frac{1}{\sqrt{1-(\frac{x}{2})^2}}\]

OpenStudy (anonymous):

the make the substitution \(u=\frac{x}{2}\) and you get it right away

OpenStudy (anonymous):

Whenever you see \(\sqrt{1-ax^2}\), \(\sqrt{1+ax^2}\), or \(\sqrt{ax^2-1}\) you should immediately think trig sub.

OpenStudy (anonymous):

or you can even make \(x=2\sin(\theta)\) and so \(dx=2\cos(\theta)d\theta\) get \[\int \frac{2\cos(\theta)d\theta}{\sqrt{4-4\sin^2(\theta)}}\]

OpenStudy (anonymous):

Then remember \[ 1-\sin^2(x) = \cos^2(x) \]\[ 1+\tan^2(x) = \sec^2(x) \]\[ \sec^2(x)-1 = \tan^2(x) \]

OpenStudy (anonymous):

the denominator is \(2\cos(\theta)\) so you are integrating \(\int 1 d\theta=\theta=\sin^{-1}(\frac{x}{2})\)